OpenStudy (anonymous):

What is the value of m in the equation 1/2m-3/4n=16 , when n=8? 20 32 44 48

OpenStudy (abdullahm):

\(\sf\LARGE\frac{1}{2}m-\frac{3}{4}n=16\) Plug in n=8 into it and solve for m \(\sf\LARGE\frac{1}{2}m-\frac{3}{4}\color{red}{(8)}=16\)

OpenStudy (anonymous):

Thanks

OpenStudy (abdullahm):

np